OpenStudy (anonymous):

Create a quadratic equation that has an axis of symmetry of x = 1 and whose graph opens down. Then find the vertex, domain, range, and x-intercepts. Show all work to receive full credit.

hero (hero):

I guess what I helped you with earlier was of no use.

OpenStudy (anonymous):

the quadratic equation needs to be in y = ax2 + bx + c form @Hero sorry

hero (hero):

Okay, well expand \((x-1)^2\) to get \(x^2 -2x + 1\)

hero (hero):

(\(x-1)^2 = x^2 - 2x + 1\)

hero (hero):

\(x^2 - 2x + 1\) is in the form \(ax^2 + bx + c\). Do you agree?

OpenStudy (anonymous):

i agree

hero (hero):

Okay, so you're good then, right?

OpenStudy (anonymous):

i put it in a graphing calculator and it doesnt open down though :(

hero (hero):

I forgot that it had to be negative: \(-(x-1)^2 = -(x^2 - 2x +1)\) All you had to do was put a negative in front of it and you get the flip side.

hero (hero):

Or you can just input \(-x^2 + 2x - 1\)

OpenStudy (anonymous):

x has to equal one though..

hero (hero):

It will still equal 1. What me to prove it to you?

OpenStudy (anonymous):

no i believe u

hero (hero):

You have a calculator bro. Just graph and you will see that x still equals 1

OpenStudy (anonymous):

i mean squared

hero (hero):

I helped you with this earlier. If you don't know how to expand binomial squares, you need to brush up on your algebra.

OpenStudy (anonymous):

i would write it out like this right −1^2+2(1)−1=0 @Hero

hero (hero):

-(1)^2 + 2(1) - 1 = 0